# Scaling Law Hypothesis for Multimodal Model

### v1

@@ Introduction
Recent work suggests a unified scaling law, where model performance is driven primarily by total compute, regardless of how it is distributed between model size and dataset size [8]. This approach introduces bits per character (BPC) as a performance metric that reflects the model’s compression efficiency. BPC has been shown to correlate linearly with model performance across various modalities (Figure 1).
Figure 1: Performance (measured by BPC) correlates with data compression ability across different text corpora [8].
@@
$\text{multimodal performance}\propto\log\left(\sum_{i}\frac{T_{i}}{C_{i}}\right)+\log P,$ (4)

### v2

@@ Introduction
Recent work suggests a unified scaling law, where model performance is driven primarily by total compute, regardless of how it is distributed between model size and dataset size [8]. This approach introduces bits per character (BPC) as a performance metric that reflects the model’s compression efficiency [9]. BPC has been shown to correlate linearly with model performance across various modalities (Figure 1).
Figure 1: Performance (measured by BPC) correlates with data compression ability across different text corpora.
@@
$\text{multimodal performance}\propto\sum_{i}\log\left(\frac{T_{i}}{C_{i}}\right)+\log P,$ (4)